OpenStudy (anonymous):

Simplify: x to the 5th power times x cubed times x to the negative 1 power

OpenStudy (carson889):

x^5 * x^3 * x^ -1 When multiplying variables with powers, you can simpply add the value of the powers together. 5 + 3 + (-1) = 7, Therefore the answer is x^7 or x to the power of seven.

OpenStudy (carson889):

If it was x to the power of five divided by x to the power of 3, then it becomes a matter of subtracting the two powers. In this example of division, the answer would be x^2 or x to the power of two.

OpenStudy (carson889):

\[\frac{ x ^{n} }{ x ^{m} } =x ^{nm}\] \[\frac{ x ^{n} }{ x ^{m} } =x ^{n-m}\] \[\left( x ^{n} \right)^{m}=x ^{n*m}\]